Operations Report – April 13th

Crew 279 Operations Report 13-04-2023

SOL: 11

Name of person filing report: Ioana Dimitrova

Non-nominal systems: robotic observatory, Opportunity

Notes on non-nominal systems:

Robotic observatory: awaiting a solution from the manufacturer.
Opportunity Rover : see rover section

Spirit rover used: yes
Hours: 216,2
Beginning charge: 100%
Ending charge: 67%
Currently charging: no
Opportunity rover used: yes
Hours: 120,5
Beginning charge: 100%
Ending charge: 89%
Currently charging: no
Curiosity rover used: yes
Hours: 229,2
Beginning charge: 100%
Ending charge: 69%
Currently charging: no
Perseverance rover used: yes
Hours: 260,9
Beginning charge: 100%
Ending charge: 87%
Currently charging: no
General notes on rovers: one of the battery wires on the driver side of Opportunity is corroded out and needs to be replaced.

Summary of Hab operations: Music influence on stress experiment was conducted.
Water (static tank): 269
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no

Summary of internet: used from 14:00 to 21:00 for experiments and reports.

Summary of suits and radios: Radios charging nominal.
Suit 2 battery was changed and suit 2 is nominal now. Faulty one is placed in RAM along with the other one.
Summary of GreenHab operations:
Heater: On
Supplemental light: Off
Harvest: none
Summary of ScienceDome operations : bacteria that were recuperated this morning were taken out of the liquid environment and put in agar medium cultures to be able to analyze them tomorrow by CFU.
Dual split: Heat or AC, Off

Summary of RAM operations: none
Summary of any observatory issues: see above for robotic observatory. Tracking error at New Mexico observatory. See Observatory report.
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site